OpenStudy (anonymous):

HELP!!! Misty deposited $6,625 into a savings account 16 years ago. The account has an interest rate of 4.1% and the balance is currently $12,724.27. How often does the interest compound?

OpenStudy (anonymous):

y=ab^x. A is the initial amount, b is the interest rate, and x is the time. y is obviously the amount that you have. Does that help at all?

OpenStudy (anonymous):

@halorazer, no that formula won't help. It's asking for how often the interest gets compoundedper year A = P (1 + r/n)^(nt) A = $12,724.27. P = $6,625 r = 0.041 t = 16 n = to be found

OpenStudy (anonymous):

I think. y=6,625(0.041/12)^16 y=6,625(0.0034)^16 y=6,625 (3.19) y=21,133.45 ? Then add that to the initial amount? (I know it's not the answer: what's left is annually, quarterly and daily.) That's how you do it, right?

OpenStudy (anonymous):

sourwing is correct, I am sorry. D; disregard what I said. Use pert.

OpenStudy (anonymous):

Oh okay. :) So.... 12,724=6,625(1+0.041/n)^16n ? But how do you work with those n's ????

OpenStudy (anonymous):

*12,724.27 sorry.

OpenStudy (anonymous):

@sourwing

OpenStudy (anonymous):

there is no algebraic way to solve for n. You will have to graphing it using a graphing calculator. n = 4

OpenStudy (anonymous):

Oh. Okay. And I plugged in all the answer choices the question gave me and I got 4! Thank you! :)